OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i A) f(x) = x^4 + 12.5x^2 - 50x - 150 B) f(x) = x^4 - 4x^3 + 15x^2 + 25x + 150 C) f(x) = x^4 - 4x^3 - 15x^2 - 25x - 150 D) f(x) = x^4 - 9x^2 - 50x - 150

OpenStudy (anonymous):

I need help fast!

OpenStudy (anonymous):

Because the polynomial has real coefficients and -1+3i is a zero of it, then -1-3i is a zero to it also ! So your polynomial is : \[\Large f(x)=(x-5)(x+3)(x-(-1+3i))(x-(-1-3i))\]

OpenStudy (anonymous):

Okay, how do I convert that to standard form?

OpenStudy (anonymous):

We have : \[\Large (x-5)(x+3)=x^2+3x-5x-15=x^2-2x-15\] and : \[\Large (x+1+3i)(x+1-3i)\\\Large=x^2+(1+3i)x+(1-3i)x+(1+3i)(1-3i) \\\Large=x^2+2x+4 \] So : \[\Large f(x)=(x^2-2x-15)(x^2+2x+4)\]

OpenStudy (anonymous):

5, -3, and -1 + 3i remamber that those have a conjugate pair. THe imaginaries (x-5)*(x+3)*(x-(-1 + 3i ))*(x+(-1 + 3i )) now expand

OpenStudy (anonymous):

OK ! It was there a mistype : \[\Large (x+1+3i)(x+1−3i)=x^2+2x+10\]

OpenStudy (anonymous):

x^4-2 x^3-(7-6 i) x^2-(16+12 i) x-(120+90 i)

OpenStudy (anonymous):

i'd go with noura